Solve the following equation:


x Ix-5I = 20x+3


List the four possible roots in increasing order and indicate if is a EXTRANEOUS root or just a ROOT.




1
Expert's answer
2021-02-24T07:38:15-0500
"x|x-5|=20x+3"

"x\geq5, |x-5|=x-5"


"x(x-5)=20x+3"

"x^2-5x-20x-3=0"

"x^2-25x-3=0"

"D=b^2-4ac=(-25)^2-4(1)(-3)=637"

"x_1=\dfrac{25-\sqrt{637}}{2(1)}=\dfrac{25-7\sqrt{13}}{2}"

"x_2=\dfrac{25+\sqrt{637}}{2(1)}=\dfrac{25+7\sqrt{13}}{2}"

"x<5, |x-5|=-(x-5)"

"-x(x-5)=20x+3"

"x^2-5x+20x+3=0"

"x^2+15x+3=0"

"D=b^2-4ac=(15)^2-4(1)(3)=213"

"x_1=\dfrac{-15-\sqrt{213}}{2(1)}=\dfrac{-15-\sqrt{213}}{2}"

"x_2=\dfrac{-15+\sqrt{213}}{2(1)}=\dfrac{-15+\sqrt{213}}{2}"


List the four possible roots in increasing order 


"\dfrac{-15-\sqrt{213}}{2}, \dfrac{-15+\sqrt{213}}{2},"

"\dfrac{25-7\sqrt{13}}{2}, \dfrac{25+7\sqrt{13}}{2}"

"x=\dfrac{-15-\sqrt{213}}{2}" just root


"x=\dfrac{-15+\sqrt{213}}{2}" just root


"x=\dfrac{25-7\sqrt{13}}{2}" extraneous root


"x=\dfrac{25+7\sqrt{13}}{2}" just root